OpenStudy (kayders1997):

Will fan and metal Find relative and extremes values of the function f(x)=x^3-6x^2+9x+1

OpenStudy (anonymous):

any choices?

OpenStudy (kayders1997):

No

OpenStudy (kayders1997):

I know you have to find the derivative

OpenStudy (anonymous):

@Brill might be able to help

OpenStudy (anonymous):

I suppose relative and extreme values in this case means local and global maxes and mins I'm not entirely sure there is a way to do this other than graphing I would recommend graphing and finding the mins and maxes

OpenStudy (kayders1997):

Thanks

OpenStudy (anonymous):

\[f \prime \left( x \right)=3x^2-12x+9\] \[f \prime \left( x \right)=0,gives~3x^2-12x+9=0,x^2-4x+3=0\] \[x^2-3x-x+3=0,x \left( x-3 \right)-1\left( x-3 \right)=0,\left( x-3 \right)\left( x-1 \right)=0,x=1,3\] \[f \prime \prime \left( x \right)=6x-12\] when x=1,find f"(x) when f"(x)<0,there is relative maxima if f "(x)>0,there is relative minima. similarly for x=3 ?

OpenStudy (kayders1997):

Okay I'm following most of this